OpenStudy (perl):

Question about code cog Latex editor. Is there a way to make a new line without losing the 'large' font. I wanted to copy paste from code cog. When I copy paste from it I get the following

OpenStudy (perl):

\[\large (2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3)=(4,6) \\ (-6,3) \to (2\cdot (-6), 2 \cdot 3)\]

OpenStudy (perl):

@e.mccormick

OpenStudy (perl):

http://prntscr.com/76yimp

OpenStudy (perl):

it seems that i have to manually place a pair of braces around \Large { } , is that the only workaround

OpenStudy (bloomlocke367):

I believe so, or you could just do the \large code again, but I think it'll be more efficient to just use { }.

OpenStudy (e.mccormick):

Yah, bloom has the right idea. Designate what you want as large using {} ``` \(\large{(2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3)=(4,6) \\ (-6,3) \to (2\cdot (-6), 2 \cdot 3)}\) ``` \(\large{(2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3)=(4,6) \\ (-6,3) \to (2\cdot (-6), 2 \cdot 3)}\)

OpenStudy (perl):

@e.mccormick Is there anyway to get a smaller negative sign, a hyphen or maybe en-dash The negative sign above looks like a subtraction sign $$\large{(2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3)=(4,6) \\ (-6,3) \to (2\cdot (-6), 2 \cdot 3)}$$ $$ \Large 3 - (\color{red}{-}6)$$ I would prefer a smaller hyphen

OpenStudy (perl):

ALso it would be nice if the codecog editor let you scroll down the preview window

OpenStudy (perl):

it says equation is too long here http://prntscr.com/77curi

OpenStudy (bloomlocke367):

I'm not sure about that..

OpenStudy (bloomlocke367):

I know a way for you to make smaller dashes, though.

OpenStudy (perl):

ok , smaller dash is good

OpenStudy (bloomlocke367):

\(\Large 3 - (\small\color{red}{-}\Large6)\) you could just change the size, like this: `\(\Large 3-(\small\color{red}-\Large6)\)`

OpenStudy (bloomlocke367):

it's not the easiest thing, but it works.

OpenStudy (perl):

thats interesting :)

OpenStudy (bloomlocke367):

do you know all of the sizes? they could be very useful.

OpenStudy (perl):

for the most part, \large, \Large, \Huge, \small, etc

OpenStudy (perl):

do you know them by heart?

OpenStudy (perl):

$$ \tiny a \Tiny a \small a \normalsize a \large a \Large a \LARGE a \Huge a$$

OpenStudy (bloomlocke367):

no, but I was just about to get a table to put here.

OpenStudy (perl):

sure you can do that

OpenStudy (perl):

also as far as my question above about the negative sign, this sounds like a good article http://tex.stackexchange.com/questions/79141/is-there-a-designated-symbol-for-the-negative-sign-in-say-16 it seems like a workaround is to use \text{-}

OpenStudy (perl):

$$\Large 3 - (\text{-}6)$$

OpenStudy (bloomlocke367):

Interesting. That's a lot easier than what I said.

OpenStudy (perl):

this works regardless of the relative sizes

OpenStudy (bloomlocke367):

\(\Huge\tt\color{lime}{I~see~that...}\) \(\Huge6+\text-2\)

OpenStudy (perl):

is there a way to prevent the latex code from going off the screen, besides for manually putting in new line. example: \( \Large { V=\int_{1}^{1+h}\pi ( \frac{1}{y})^2 ~dy =\int_{1}^{1+h} \frac{\pi}{y^2} ~dy =\int_{1}^{1+h} {\pi}{y^{-2}} ~dy = [\pi \cdot \frac{y^{-1}}{-1}] _{1}^{1+h} = -\pi h^{-1} - (-\pi \cdot 1 ) = \pi - \frac{\pi}{h} }\)

OpenStudy (bloomlocke367):

you can make it smaller... but if it's too small, it can be difficult to read.

OpenStudy (e.mccormick):

From what I have seen, MathJax does not do automatic line wrapping. Unicode U+2013 is the endash, so let's see... \(\unicode{x2013}\) ` \(\unicode{x2013}\)` \(6+\unicode{x2013}2\) \(\large 6+\unicode{x2013}2\) \(\Large 6+\unicode{x2013}2\) \(\LARGE 6+\unicode{x2013}2\) \(\huge 6+\unicode{x2013}2\) \(\Huge 6+\unicode{x2013}2\)